# SigFormer: Sparse Signal-Guided Transformer for Multi-Modal Human Action Segmentation

### v1

@@ 1. INTRODUCTION
To tackle the challenge of multi-modal action segmentation utilizing sparse and dense signals, we propose a Sparse signal-guided Transformer (SigFormer). Our objective is to explore an effective method to utilize multi-modal signals to attain accurate representations of fine-grained actions. In SigFormer, we propose a sparse guided cross-modal module designed to integrate multi-modal signals with various information densities, aiming to achieve a comprehensive representation of human actions. It consists of two modules: Sparse Guided Fusion (SGF) module and Motion-Spatial Attention Fusion (MSAF) module. The SGF module is responsible for fusing sparse state signals from IoT-enabled devices with dense signals. Within the SGF module, we utilize masked attention to restrict the attention to the localized feature fusion within temporal segments where sparse signals are valid. Compared to other fusion methods and the typical cross-attention used in a standard Transformer, our SGF module with masked attention leads to superior performance. The MSAF module is tasked with encoding the interaction context between motion and spatial information that has been precisely guided by sparse signals. Furthermore, we emphasize the awareness of temporal action boundary at two stages. First, we introduce an intermediate bottleneck module to acquire task-specific aligned features for action segmentation. In this module, instead of concentrating solely on action categories (Uchiyama, 2023; Carreira and Zisserman, 2017; Li et al., 2022), we enhance the distinctive features of each modality, encompassing both action category and boundary information by incorporating inner losses into the end-to-end network. This ensures that the features of multiple modalities are aligned within a consistent semantic space, thereby enabling the extraction of valuable information related to action segmentation. Additionally, we incorporate two mutual interactive branches after feature fusion, which explicitly model the multi-stage interaction between the category branch and the boundary branch to further enhance the boundary awareness of the overall network. By placing great emphasis on temporal action boundaries, we effectively alleviate over-segmentation errors. Evaluation results of SigFormer on a large-scale multi-modal benchmark from real industrial environments demonstrate that it can achieve better results compared to state-of-the-art methods.
@@
  • To emphasize the temporal action boundary, we introduce an intermediate bottleneck module to jointly learn both category and boundary of each dense modality through the inner loss functions. Furthermore, we facilitate multi-stage interactions between action category and temporal boundary in mutual interactive branches to further augment the overall network’s boundary awareness.
@@ 2. RELATED WORK
Multi-Modality Action Recognition. Multi-modality action recognition techniques have gained significant attention due to the complementary information offered by different modalities, leading to improved classification performance through the fusion of multiple feature streams. Previous approaches mainly utilized decision-level fusion strategies (Cai et al., 2021; Hu et al., 2017; Wu et al., 2016; Zhang et al., 2018; Zhao et al., 2017), such as score fusion, or relied on simplistic feature-level fusion techniques. For example, Liu and Yuan (Liu and Yuan, 2018) concatenated skeleton and pose heatmaps features. Noori et al. (Noori et al., 2020) explored several methods of fusion for multi-representations of data from sensors and proposed a CNN-based general architecture. Recent methods have focused on designing correlation modules for modality fusion (Li et al., 2020; Joze et al., 2020; Cai et al., 2021; Zhang et al., 2022). Li et al. (Li et al., 2020) developed bilinear pooling to enable the skeleton feature to guide on RGB feature at the semantic feature level. Joze et al. (Joze et al., 2020) introduced the fusion of multi-modality features across different spatial dimensions, enabling the re-calibration of channel-wise features within each CNN stream utilizing the knowledge from multiple modalities. Zhang et al. (Zhang et al., 2022) turned the video action segmentation and recognition problem into node classification of graphs and utilized GNNs to learn multi-modal feature fusion.
@@
Recently, there has been a shift in research focus towards Transformers (Nagrani et al., 2021; Wang et al., 2022; Xu et al., 2022; Li et al., 2022; van Amsterdam et al., 2022, 2023), which are versatile perceptual models capable of effectively handling diverse data sources with minimal modifications to the model structure. Although the attention operation in Transformers provides a straightforward approach to integrating various signals, it fails to consider information redundancy and exhibits limited scalability when applied to longer temporal sequences. To address these challenges, Nagrani et al. (Nagrani et al., 2021) proposed an attention bottleneck to control the flow of attention between tokens from different data sources, ensuring that each modality shares only essential information with the others. Van et al. (van Amsterdam et al., 2023) explicitly distinguished between modality-shared and private feature representations within a long-range action segmentation model, leveraging the bottleneck mechanism to maintain the separation of features in subsequent temporal processing layers. However, these methods do not ensure the quality of multi-modal features before fusion, which might result in the integration of irrelevant information and the introduction of interference. In this paper, we propose an intermediate bottleneck module to regulate the feature extraction process of diverse data and employ accurate sparse signals to guide dense signals.
@@ 3. METHODOLOGY
This section provides a comprehensive description of SigFormer. SigFormer takes dense signals, including IMU data (I), body key points (K), body bounding box points (B), as well as sparse state signals of IoT-enabled devices (D) as input and outputs frame-level refined class predictions. We commence with an overview of SigFormer, subsequently delving into an intricate presentation of its four constituent modules.
@@ 3.1. Overview
Figure 1. Overview of the pipeline for SigFormer. (a) Feature extraction from each signal. (b) The intermediate bottleneck module operates by taking $F_{i}$ ($i$ $\in M,K,B$) as input, reducing its feature dimension, calculating the inner category loss $L_{c_{inner}}$, and the inner boundary loss $L_{b_{inner}}$. Subsequently, it restores the feature dimension to obtain $F_{o}$ ($o$ $\in M,K,B$), which possesses an enhanced awareness of both category and boundary. (c) The sparse guided cross-modal module combines multi-modal signals features with different information densities, $F_{M}$, $F_{D}$, and $F_{S}$ (obtained by adding $F_{K}$ and $F_{B}$), to generate a comprehensive action representation $F_{C}$. This module comprises a sparse guided fusion (SGF) module and a motion-spatial attention fusion (MSAF) module. (d) Mutual interactive branches employ Multi-Head Cross Attention (MHCA) to model information interaction between the class branch and boundary branch. The predictions of the class branch are refined by the boundaries (filtered from the predictions of the boundary branch) to obtain frame-level refined class predictions.
@@
Figure 1. Overview of the pipeline for SigFormer. (a) Feature extraction from each signal. (b) The intermediate bottleneck module operates by taking $F_{i}$ ($i$ $\in M,K,B$) as input, reducing its feature dimension, calculating the inner category loss $L_{c_{inner}}$, and the inner boundary loss $L_{b_{inner}}$. Subsequently, it restores the feature dimension to obtain $F_{o}$ ($o$ $\in M,K,B$), which possesses an enhanced awareness of both category and boundary. (c) The sparse guided cross-modal module combines multi-modal signals features with different information densities, $F_{M}$, $F_{D}$, and $F_{S}$ (obtained by adding $F_{K}$ and $F_{B}$), to generate a comprehensive action representation $F_{C}$. This module comprises a sparse guided fusion (SGF) module and a motion-spatial attention fusion (MSAF) module. (d) Mutual interactive branches employ Multi-Head Cross Attention (MHCA) to model information interaction between the class branch and boundary branch. The predictions of the class branch are refined by the boundaries (filtered from the predictions of the boundary branch) to obtain frame-level refined class predictions.
@@ 3.3. Intermediate Bottleneck Module
Most approaches (Farha and Gall, 2019; Ishikawa et al., 2021; Yi et al., 2021; Li et al., 2022) for action segmentation typically rely on pre-trained feature extractors (Carreira and Zisserman, 2017; Li et al., 2022) for data processing and feature extraction, which are subsequently followed by feature fusion. However, these methods lack emphasis on capturing action boundary information within the data. Consequently, the resulting features inadequately localize the temporal boundaries between actions, thereby falling short of meeting the specific requirements of action segmentation. In contrast, we introduce a bottleneck module subsequent to the feature extraction of each modality, thus imposing additional constraints. Through the enhancement of awareness regarding action class and temporal boundary, this module ensures that exclusively pertinent information essential for action segmentation is extracted from each modality. As illustrated in Fig. 1 (b), the architecture of the intermediate bottleneck module contains two parts: the inner class branch and the inner boundary branch.
@@
where $F_{class}\in\mathbb{R}^{N\times T\times C}$ is the class output feature, $C$ represents the number of action classes ($C$ $E$), $P_{class}$ is the predicted result of the inner class, from which the inner class loss $\mathcal{L}_{c_{inner}}$ is calculated. The feature $F_{class}\in\mathbb{R}^{N\times T\times C}$ is then fed into another MLP layer $f_{p}^{{}^{\prime}}(\cdot)$, which adjustments the feature dimension back. Additionally, to prevent the potential loss of feature information resulting from intermediate dimensional reduction, we incorporate the input of the intermediate bottleneck module $F_{i}\in\mathbb{R}^{N\times T\times E}$ into the output of $f_{p}^{{}^{\prime}}(\cdot)$ via a residual connection (He et al., 2016), ensuring the preservation of the original feature information:
@@
where $P_{boundary}\in\mathbb{R}^{N\times T\times 1}$ is employed to calculate the inner boundary loss $\mathcal{L}_{b_{inner}}$.
In conclusion, the intermediate bottleneck module further constrains the feature extraction stage by calculating the inner class loss and the inner boundary loss. It introduces class and boundary awareness ability upfront for each modality and ensures the provision of precise and valuable information for modality fusion.
@@ 3.4. Sparse Guided Cross-Modal Module
Here, $d$ is a normalization constant, $\ast$ is element-wise product, guided enhanced features provide keys and values for MHMA, and sparse guidance provides queries. Moreover, the attention mask $\mathcal{M}$ at feature location is
@@ 3.5. Mutual Interactive Branches
For the problem of over-segmentation, most previous work utilized multi-stage refinement (Farha and Gall, 2019; Li et al., 2023; Yi et al., 2021) and proposed boundary-aware networks (Wang et al., 2020a; Ishikawa et al., 2021), etc. However, these studies disregard the intrinsic interplay between action class and temporal boundary, making it challenging to achieve proper trade-offs and balance during refinement. Instead, we explicitly model the interrelationship between action class and boundary within the mutual interactive branches.
@@ 3.6. Loss Function
(12) $\mathcal{L}_{c}=\frac{1}{T}\sum_{t}{-\log(y_{t,c})}$
where $y_{t,c}$ is the frame-wise prediction action probability for class $c$ at time $t$.
@@
where $y_{t}$ and $p_{t}$ are the ground truth and the action boundary probability for frame $t$, respectively.
Total Loss. The SigFormer generates the inner loss ( inner class loss $\mathcal{L}_{c_{inner}}$ and inner boundary loss $\mathcal{L}_{b_{inner}}$) through the intermediate bottleneck module and the final loss ( final class loss $\mathcal{L}_{c_{final}}$ and final boundary loss $\mathcal{L}_{b_{final}}$) through the mutual interactive branches. They are combined linearly as follows:
(14) $\mathcal{L}_{Total}=\lambda_{1}\mathcal{L}_{c_{inner}}+(1-\lambda_{1})\mathcal{L}_{c_{final}}+\lambda_{2}(\mathcal{L}_{b_{inner}}+\mathcal{L}_{b_{final}})$
where $\mathcal{L}_{Total}$ is the total loss of the SigFormer, $\lambda_{1}$ is utilized to balance the inner loss and final loss, that is set to 0.5 in our experiments, and $\lambda_{2}$ is employed to balance the class branch and boundary branch, which is set to 0.2 in our experiments.
@@ 4.1. Dataset
OpenPack (Yoshimura et al., 2022b) is a multi-modal large-scale dataset that is specifically designed for action segmentation in industrial domains. It contains 53.8 hours of multi-modal sensor data collected from 16 different subjects with varying levels of packaging work experience, including key points, depth images, sensory data from body-worn inertial measurement units, LiDAR point clouds, and readings from IoT-enabled devices (e.g., handheld barcode scanners utilized in work procedures). It consists of 20129 instances of activities (operations) and 52529 instances of actions with 9 different modalities. Since some modalities are incomplete, we select four of them that are complete and crucial in our experiments: IMU data, human key points, bounding box points, and state signals of IoT-enabled devices.

### v2

# SigFormer: Sparse Signal-Guided Transformer for Multi-Modal Action Segmentation

@@ 1. INTRODUCTION
To tackle the challenge of multi-modal action segmentation utilizing sparse and dense signals, we propose a Sparse signal-guided Transformer (SigFormer). Our objective is to explore an effective method to utilize multi-modal signals to attain accurate representations of fine-grained actions. In SigFormer, we propose a sparse guided cross-modal module designed to integrate multi-modal signals with various information densities, aiming to achieve a comprehensive representation of human actions. It consists of two modules: Sparse Guided Fusion (SGF) module and Motion-Spatial Attention Fusion (MSAF) module. The SGF module is responsible for fusing sparse state signals from IoT-enabled devices with dense signals. Within the SGF module, we utilize masked attention to restrict the attention to the localized feature fusion within temporal segments where sparse signals are valid. Compared to other fusion methods and the typical cross-attention used in a standard Transformer, our SGF module with masked attention leads to superior performance. The MSAF module is tasked with encoding the interaction context between motion and spatial information that has been precisely guided by sparse signals. Furthermore, we emphasize the awareness of temporal action boundaries at two stages. First, we introduce an intermediate bottleneck module to acquire task-specific aligned features for action segmentation. In this module, instead of concentrating solely on action categories (Uchiyama, 2023; Carreira and Zisserman, 2017; Li et al., 2022), we enhance the distinctive features of each modality, encompassing both action category and boundary information by incorporating inner losses into the end-to-end network. This ensures that the features of multiple modalities are aligned within a consistent semantic space, thereby enabling the extraction of valuable information related to action segmentation. Additionally, we incorporate two mutual interactive branches after feature fusion, which explicitly model the multi-stage interaction between the category branch and the boundary branch to further enhance the boundary awareness of the overall network. By placing great emphasis on temporal action boundaries, we effectively alleviate over-segmentation errors. Evaluation results of SigFormer on a large-scale multi-modal benchmark from real industrial environments demonstrate that it can achieve better results compared to state-of-the-art methods.
@@
  • To emphasize the temporal action boundary, we introduce an intermediate bottleneck module to jointly learn both the category and boundary of each dense modality through the inner loss functions. Furthermore, we facilitate multi-stage interactions between action category and temporal boundary in mutual interactive branches to further augment the overall network’s boundary awareness.
@@ 2. RELATED WORK
Multi-Modality Action Recognition. Multi-modality action recognition techniques have gained significant attention due to the complementary information offered by different modalities, leading to improved classification performance through the fusion of multiple feature streams. Previous approaches mainly utilized decision-level fusion strategies (Cai et al., 2021; Hu et al., 2017; Wu et al., 2016; Zhang et al., 2018; Zhao et al., 2017), such as score fusion, or relied on simplistic feature-level fusion techniques. For example, Liu and Yuan (Liu and Yuan, 2018) concatenated skeleton and pose heatmap features. Noori et al. (Noori et al., 2020) explored several methods of fusion for multi-representations of data from sensors and proposed a CNN-based general architecture. Recent methods have focused on designing correlation modules for modality fusion (Li et al., 2020; Joze et al., 2020; Cai et al., 2021; Zhang et al., 2022). Li et al. (Li et al., 2020) developed bilinear pooling to enable the skeleton feature to guide on RGB feature at the semantic feature level. Joze et al. (Joze et al., 2020) introduced the fusion of multi-modality features across different spatial dimensions, enabling the re-calibration of channel-wise features within each CNN stream utilizing the knowledge from multiple modalities. Zhang et al. (Zhang et al., 2022) turned the video action segmentation and recognition problem into node classification of graphs and utilized GNNs to learn multi-modal feature fusion.
Recently, there has been a shift in research focus towards Transformers (Nagrani et al., 2021; Wu et al., 2021; Liu et al., 2021; Yao et al., 2022; Li et al., 2023b; Mao et al., 2022; Wang et al., 2022; Xu et al., 2022a; Li et al., 2022; van Amsterdam et al., 2022; Yao et al., 2023; van Amsterdam et al., 2023; Zhao et al., 2022; Xu et al., 2022b), which are versatile perceptual models capable of effectively handling diverse data sources with minimal modifications to the model structure. For example, Wu et al. (Wu et al., 2021) introduced convolutions into Vision Transformer (ViT) to improve ViT in performance and efficiency. Liu et al. (Liu et al., 2021) proposed the Swin Transformer, a hierarchical Transformer with Shifted windows. Yao et al. (Yao et al., 2022) constructed a Wavelet ViT that enables self-attention learning with lossless down-sampling over keys/values. Li et al. (Li et al., 2023b) designed a Contextual Transformer to fully capitalize on the contextual information among input keys to guide the learning of dynamic attention matrices. Mao et al. (Mao et al., 2022) proposed Robust ViT which has advanced robustness and generalization ability. To explore the holistic interactions among all patches, Yao et al. (Yao et al., 2023) proposed Dual-ViT that elegantly exploits the global semantics for self-attention learning.
@@
Although the attention operation in Transformers provides a straightforward approach to integrating various signals, it fails to consider information redundancy and exhibits limited scalability when applied to longer temporal sequences. To address these challenges, Nagrani et al. (Nagrani et al., 2021) proposed an attention bottleneck to control the flow of attention between tokens from different data sources, ensuring that each modality shares only essential information with the others. Van et al. (van Amsterdam et al., 2023) explicitly distinguished between modality-shared and private feature representations within a long-range action segmentation model, leveraging the bottleneck mechanism to maintain the separation of features in subsequent temporal processing layers. However, these methods do not ensure the quality of multi-modal features before fusion, which might result in the integration of irrelevant information and the introduction of interference. In this paper, we propose an intermediate bottleneck module to regulate the feature extraction process of diverse data and employ accurate sparse signals to guide dense signals.
@@ 3. METHODOLOGY
This section provides a comprehensive description of SigFormer. SigFormer takes dense signals, including IMU data (I), body key points (K), body bounding box points (B), as well as sparse state signals of IoT-enabled devices (D), as input, and produces frame-level refined class predictions as output. We commence with an overview of SigFormer, subsequently delving into an intricate presentation of its four constituent modules.
@@ 3.1. Overview
Figure 1. Overview of the pipeline for SigFormer. (a) Feature extraction from each signal. (b) The intermediate bottleneck module operates by taking $F_{i}$ ($i$ $\in M,K,B$) as input, reducing its feature dimension, calculating the inner class loss $\mathcal{L}_{c}^{inner}$, and the inner boundary loss $\mathcal{L}_{b}^{inner}$. Subsequently, it restores the feature dimension to obtain $F_{o}$ ($o$ $\in M,K,B$), which possesses an enhanced awareness of both category and boundary. (c) The sparse guided cross-modal module combines features from multi-modal signals with different information densities, $F_{M}$, $F_{D}$, and $F_{S}$ (obtained by adding $F_{K}$ and $F_{B}$), to generate a comprehensive action representation $F_{C}$. This module comprises a sparse guided fusion (SGF) module and a motion-spatial attention fusion (MSAF) module. (d) Mutual interactive branches employ Multi-Head Cross Attention (MHCA) to model information interaction between the class branch and boundary branch. The predictions of the class branch are refined by the boundaries (filtered from the predictions of the boundary branch) to obtain frame-level refined class predictions.
[…]
@@ 3.3. Intermediate Bottleneck Module
Most approaches (Farha and Gall, 2019; Ishikawa et al., 2021; Yi et al., 2021; Li et al., 2022) for action segmentation typically rely on pre-trained feature extractors (Carreira and Zisserman, 2017; Li et al., 2022) for data processing and feature extraction, which are subsequently followed by feature fusion. However, these methods lack emphasis on capturing action boundary information within the data. Consequently, the resulting features inadequately localize the temporal boundaries between actions, thereby falling short of meeting the specific requirements of action segmentation. In contrast, we introduce a bottleneck module after the feature extraction of each modality, thus imposing additional constraints. Through the enhancement of awareness regarding action class and temporal boundary, this module ensures that exclusively pertinent information essential for action segmentation is extracted from each modality. As illustrated in Fig. 1 (b), the architecture of the intermediate bottleneck module contains two parts: the inner class branch and the inner boundary branch.
@@
where $F_{class}\in\mathbb{R}^{N\times T\times C}$ is the class output feature, $C$ represents the number of action classes ($C$ $E$), $P_{class}$ is the predicted result of the inner class, from which the inner class loss $\mathcal{L}_{c}^{inner}$ is calculated. The feature $F_{class}\in\mathbb{R}^{N\times T\times C}$ is then fed into another MLP layer $f_{p}^{{}^{\prime}}(\cdot)$, which adjustments the feature dimension back. Additionally, to prevent the potential loss of feature information resulting from intermediate dimensional reduction, we incorporate the input of the intermediate bottleneck module $F_{i}\in\mathbb{R}^{N\times T\times E}$ into the output of $f_{p}^{{}^{\prime}}(\cdot)$ via a residual connection (He et al., 2016), ensuring the preservation of the original feature information:
@@
where $P_{boundary}\in\mathbb{R}^{N\times T\times 1}$ is employed to calculate the inner boundary loss $\mathcal{L}_{b}^{inner}$.
In conclusion, the intermediate bottleneck module further constrains the feature extraction stage by calculating the inner class loss and the inner boundary loss. It introduces class and boundary awareness abilities upfront for each modality, ensuring the provision of precise and valuable information for modality fusion.
@@ 3.4. Sparse Guided Cross-Modal Module
Here, $d$ represents a normalization constant, $\ast$ denotes element-wise product, guided enhanced features provide keys and values for MHMA, and sparse guidance provides queries. Moreover, the attention mask $\mathcal{M}$ at feature location is
@@ 3.5. Mutual Interactive Branches
For the problem of over-segmentation, most previous work utilized multi-stage refinement (Farha and Gall, 2019; Li et al., 2023a; Yi et al., 2021) and proposed boundary-aware networks (Wang et al., 2020a; Ishikawa et al., 2021), etc. However, these studies disregard the intrinsic interplay between action class and temporal boundary, making it challenging to achieve proper trade-offs and balance during refinement. Instead, we explicitly model the interrelationship between action class and boundary within the mutual interactive branches.
@@ 3.6. Loss Function
(12) $\mathcal{L}_{c}=\frac{1}{T}\sum_{t}^{T}{-\log(p_{t,c})}$
where $y_{t,c}$ is the frame-wise prediction action probability for class $c$ at time $t$. Features from IMU data, body key points, and body bounding box points are processed separately through the intermediate bottleneck module. In the inner class branch, we calculate the class loss for each feature according to Equation 12. Summing these three class losses yields the inner class loss, denoted as $\mathcal{L}_{c}^{inner}$. Furthermore, the mutual interactive branches eventually yield frame-level refined class predictions, and the final class loss is computed utilizing Equation 12, denoted as $\mathcal{L}_{c}^{final}$.
@@
where $y_{t}$ and $p_{t}$ are the ground truth and the action boundary probability for time $t$, respectively. Similar to the inner class branch, the inner boundary branch computes boundary loss for each feature utilizing Equation 13, summed to derive the inner boundary loss, denoted as $\mathcal{L}_{b}^{inner}$. In addition, the output of the boundary branch within the mutual interactive branches is utilized to calculate the final boundary loss with Equation 13, denoted as $\mathcal{L}_{b}^{final}$.
Total Loss. Consider that action segmentation entails not only categorizing action classes but also accurately identifying temporal boundaries between actions. Therefore, we aim to enhance each modal feature’s awareness of action category and temporal action boundary by constraining the feature extraction process utilizing $\mathcal{L}_{c}^{inner}$ and $\mathcal{L}_{b}^{inner}$. This ensures relevant information for action segmentation is extracted. Additionally, we aim to improve the overall model’s ability to predict temporal action boundaries with $\mathcal{L}_{b}^{final}$, ensuring that the refinement phase results in smooth and stable action class predictions. $\mathcal{L}_{c}^{final}$ is utilized to optimize the final segmentation results of the model. The four losses are combined linearly as follows:
(14) $\mathcal{L}_{Total}=\lambda_{1}\mathcal{L}_{c}^{inner}+(1-\lambda_{1})\mathcal{L}_{c}^{final}+\lambda_{2}(\mathcal{L}_{b}^{inner}+\mathcal{L}_{b}^{final})$
where $\mathcal{L}_{Total}$ is the total loss of the SigFormer, $\lambda_{1}$ is utilized to balance the inner loss and final loss, that is set to 0.5 in our experiments, and $\lambda_{2}$ is employed to balance the class branch and boundary branch, which is set to 0.2 in our experiments. During training, we optimize the $\mathcal{L}_{Total}$ to enhance the overall performance of the model.
@@ 4.1. Dataset
OpenPack (Yoshimura et al., 2022b) is a multi-modal large-scale dataset that is specifically designed for action segmentation in industrial domains. It contains 53.8 hours of multi-modal sensor data collected from 16 different subjects with varying levels of packaging work experience, including key points, depth images, sensory data from body-worn inertial measurement units, LiDAR point clouds, and readings from IoT-enabled devices (e.g., handheld barcode scanners utilized in work procedures). It consists of 20,129 instances of activities (operations) and 52,529 instances of actions with 9 different modalities. Since some modalities are incomplete, we select four of them that are complete and crucial in our experiments: IMU data, human key points, bounding box points, and state signals of IoT-enabled devices.